MICROWAVE SPECTROSCOPY OF THE HEAVY-ATOM CARBENE ANALOG HGeBr

LU KANG, Department of Natural Sciences, Union College, Barbourville, KY 40906; FUMIE X. SUNAHORI, Department of Chemistry, University of Kentucky, Lexington, KY 40506; AN-DREA J. MINEI, Department of Chemistry, Wesleyan University, Middletown, CT 06459; DEN-NIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506; and \underline{STEWART E. NOVICK}, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

Eight isotopologues each of HGeBr and of DGeBr have been studied in natural abundance by pulsed-jet Fourier transform microwave (FTMW) spectroscopy. Neon was passed over dry ice cooled H_{3}GeBr or D_{3}GeBr and introduced into the pulsed valve of the FTMW spectrometer. The HGeBr and the DGeBr were produced in-situ with a 1000 V discharge of the gas mixture immediately following the nozzle opening. Only a -type transitions can be observed from 5 - 25 GHz; K_{a} = 0 transitions for the HGeBr and K_{a} = 0 and 1 transitions for the DGeBr isotopologues. The structure and bonding of the molecule will be discussed.